OpenStudy (anonymous):

A girl's feet are negative 4 over 5 yards from the surface of a pool. A boy's feet are negative 2 over 5 yards from the surface of the pool. Determine whose feet are closer to the surface.

OpenStudy (arindameducationusc):

one minute, i am helping someone else.... one minute

OpenStudy (anonymous):

tell me - which fraction is bigger \[-\frac{ 2 }{ 5 } \] or \[-\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

no - since the fractions are negative

OpenStudy (anonymous):

that is like saying -2 is bigger than -1

OpenStudy (anonymous):

the bigger fraction is incorrect

OpenStudy (anonymous):

so the answer would be -2/5 or the boy's feet

OpenStudy (anonymous):

The girl's feet are closer to the surface of the pool because negative 4 over 5 > negative 2 over 5. The boy's feet are closer to the surface of the pool because negative 4 over 5 > negative 2 over 5. The girl's feet are closer to the surface of the pool because negative 4 over 5 < negative 2 over 5. The boy's feet are closer to the surface of the pool because negative 4 over 5 < negative 2 over 5.

OpenStudy (texaschic101):

-2/5 = - 0.4 -4/5 = - 0.8 we need to find the smallest distance from the pool. With positive numbers, the bigger number is greater, but with negative numbers, the smaller number is greater. Therefore, -0.8 (or -4/5) is greater then -0.4 (or -2/5). Boys feet are closer because -4/5 > -2/5

OpenStudy (anonymous):

you are wrong